# Comprehensive Review of Deep Unfolding Techniques for Next-Generation Wireless Communication Systems

### v2

@@ VII-A MIMO ISAC
mmWave MIMO is a key technology in 6G wireless communication, offering high transmission capacity and sub-meter user localization. Its large-scale antenna arrays enable precise angular resolution. Traditional localization methods estimate position-related parameters before reconstructing the user’s location. To improve this, Fan et al. proposed a fast direct localization method in [89]. They use an unfolded ADMM solver with low complexity for direct localization. The trainable parameters include $\rho$ (penalty parameter), $w$ (weight coefficient for channel gain between the base station and user), and $\tau_{1}$, $\tau_{2}$ (two positive proximal parameters).
VII-C Modern Multicarrier Modulation for ISAC
5G and beyond 5G wireless communication systems operate at carrier frequencies in the microwave regime, where radar applications are dominant. This has led to the development of joint waveforms for sensing and communication to address limited spectral resources [210]. OFDM-based radars in LTE and 5G NR mobile networks’ BS are studied for radar and sensing in [211], [212], [213]. OTFS modulation for joint radar parameter estimation and communication is explored in [214], [215], [216].
An ISAC signal processing problem was proposed in [102] to process pilot and data signals together. The optimization problem requires the simultaneous recovery of CSI, transmitted data symbols, and target information, making it non-convex. To solve this, the authors used an iterative optimization algorithm that alternates between passive sensing and communication detection. Jingcai et al. proposed a deep unfolded network for target parameter estimation in MIMO-OFDM-based ISAC systems [103]. Their method leverages the sparsity of target parameters, using a compressed sensing approach. The problem is reformulated as a least absolute shrinkage and selection operator (LASSO) problem. The fast ISTA algorithm is then unfolded, with learnable parameters enabling high-precision solutions with low computational complexity.
The problem of estimating parameters of moving targets in high-speed scenarios using OTFS modulation for ISAC systems is explored in [104]. The target parameter estimation is reformulated as a compressed sensing problem. For the deep unfolded network, ADMM iterations are unfolded into the network’s layers.
VII-D IRS-Assisted ISAC
Liu et al. explored beamforming design for an IRS-assisted ISAC system with bounded angle uncertainty of sensing targets and statistical CSI uncertainty related to user equipment in [117]. They formulated a bi-level optimization problem, unfolding the projection method and Riemannian MO approach. These methods handle the maximum transmit power and unit-modulus constraints. The bi-level problem consists of two parts: the lower-level problems focus on identifying the worst-case CSI and target sensing angles, while the upper-level problem addresses beamforming design and optimization of Lagrange multipliers based on the given CSI and target angles. A double-loop deep unfolding NN is applied, unfolding the iterative gradient descent into a multi-layer structure, with each layer containing an upper-level and a lower-level loop. Trainable step sizes and offset parameters are introduced as network parameters.
VIII Decoding in Error Correcting Codes
Error control codes are essential for reliable and efficient communication. During data transmission, bits may flip due to channel noise, causing errors. Error correction coding helps detect and correct these errors at the receiver. Claude Shannon revolutionized channel coding with his landmark paper [217].
Mobile communication has evolved through several generations with significant technological advancements. In the first generation (1G), communication was primarily analog, leaving no need for channel coding. With the second generation (2G), convolutional and block codes were introduced to compress voice signals and convert them into digital form. Turbo codes were adopted as standard coding schemes in the third (3G) and fourth (4G) generations, offering near-Shannon limit performance with moderate decoding complexity [218]. Low-density parity-check (LDPC) codes, first introduced in the 1960s, were later rediscovered and extensively studied [219], [220], [221]. They are widely used in 4G, 5G, and beyond. LDPC codes are represented using a bipartite graph, placing them in the “codes on graphs” category, designed for capacity-approaching codes with iterative decoding algorithms. Since LDPC codes use sparse graphs, the number of edges grows linearly with block length, making them low-density and less complex [222]. Polar codes, introduced by Erdal Arıkan in 2009, are a modern class of error-correcting codes designed to achieve the capacity of symmetric binary-input discrete memoryless channels [223]. They are based on channel polarization and are known for their capacity-achieving property, meaning they can approach the Shannon limit as block length increases.
BP is widely used to decode modern error correction codes [224]. This decoder is constructed from the Tanner graph, introduced by Michael Tanner in his landmark paper [225]. He developed the min-sum and max-sum algorithms, which are derived from the generic sum-product decoding [226]. Polar codes use successive cancellation decoding, where received bits are decoded sequentially by leveraging frozen bits and previously decoded bits. More advanced techniques, such as successive list decoding assisted by cyclic redundancy check (CRC), further improve performance [227].
In practical wireless communication, channels often deviate from mathematical models or are not precisely known at the decoder. In such cases, a DL-based decoder offers a robust and adaptive solution. Additionally, model-driven DL is well-suited for providing an interpretable analysis of iterative decoding algorithms. Wei et al. [228] proposed an unfolded ADMM-based penalized algorithm for binary linear codes. Their approach introduces two networks that transform the ADMM penalty parameters $\{\mu,\alpha\}$ into iteration-dependent parameters, along with an adjustable penalty function. Each stage of the learnable ADMM decoding network (LADN) corresponds to one ADMM iteration, where iteration-dependent parameters enhance convergence and reduce the number of iterations. The key idea is that increasing the number of learnable parameters improves generalization capability. For the second network, a piecewise linear penalty function with adjustable slopes is designed to suppress pseudo codewords effectively, improving decoding performance beyond conventional $\ell_{1}$ and $\ell_{2}$ norm functions. Parameter updates are trained using backpropagation via the Adam optimizer. A customized loss function based on MSE ensures effectiveness at every iteration by considering outputs from all layers. The authors claim that their decoder outperforms the conventional ADMM decoder. Additionally, decoding irregular binary LDPC codes via unfolded ADMM is explored in [229].
In [48], the authors propose the TurboNet architecture, which combines DL with the max-log MAP algorithm. While most existing model-driven decoding algorithms rely on BP and are iterative, the authors enhance the traditional MAP algorithm by unfolding it, with each iteration represented as a DNN decoding unit. As a result, the soft-input soft-output decoder, which traditionally uses the max-log MAP algorithm and extrinsic LLRs, is replaced by two subnets based on a neural log-MAP algorithm for each iteration.
In [230], the authors tackle the decoding challenges of marker codes in insertion and deletion channels, with a focus on overcoming CSI uncertainty. To improve detection performance, they propose DL techniques, implementing both RNNs and a deep unfolded network. The forward-backward algorithm is unfolded into multiple layers within the network to enhance the overall decoding process.
Gao et al. [231] have unfolded the factor graph of the BP decoder using a one-bit quantizer for polar codes. In [232], a double-loop iterative decoding algorithm for LDPC codes is unfolded within a penalty dual decomposition framework. Additionally, by unfolding the iterative decoding process between check nodes and variable nodes into a feed-forward propagation network, a deep unfolded normalized min-sum LDPC decoder is derived in [233].
@@ VIII Decoding in Error Correcting Codes
Tanner-graph-based transmission schemes for massive multiple access, incorporating NOMA and grant-based random access, are discussed in [234]. To enhance decoding efficiency, a message-passing decoder is unfolded for both transmission schemes, allowing the decoder to improve as more information about the transmitted data is learned during the training epochs. In [235], theoretical results for the unfolded BP decoder are presented. The authors define the generalization gap of a decoder as the difference between the empirical and expected bit-error rates. They provide a set of theoretical results that bound this gap and examine its dependence on decoder complexity, which is influenced by code parameters (such as block length, message length, and variable/check node degrees), the number of decoding iterations, and the size of the training dataset.
IX Power allocation
In a multi-antenna wireless communication system with multiple users and limited resources, ensuring reliable transmission of massive amounts of data presents significant challenges. One of the primary constraints is the power budget, which must be allocated efficiently among data streams to maintain a high quality of service. The water-filling algorithm is a widely recognized technique for optimal power allocation across subchannels, aiming to maximize the sum capacity or sum rate. Since power allocation directly impacts channel capacity, it is formulated as an optimization problem, where maximizing the capacity rate leads to improved system performance. The power allocation problem can thus be formulated as an optimization problem as shown in (6) below:
$\max_{\left\{P_{i}\right\}}$ $\frac{\sum_{i=1}^{K}w_{i}R_{i}(P_{i})}{\sum_{i=1}^{K}P_{i}}$ (6)
s.t. $R_{i}(P_{i})\geq R_{\min,i},\quad\forall i,$
$\sum_{i=1}^{K}P_{i}\leq P_{\text{total}},$
$P_{i}\geq 0,\quad\forall i.$
where $K$ is the number of users, $w_{i}$ is the weight associated to user $i$ decided by user priority or fairness, $R_{i}(P_{i})$ is the rate achieved by user $i$ given the power $P_{i}$ allocated to them, $R_{\text{mini},i}$ is the minimum required rate for fairness and $P_{\text{total}}$ is the sum of all users’ power values.
@@ IX Power allocation
Several classical approaches, such as the Dinkelbach algorithm, Lagrange dual decomposition, interference pricing, successive convex approximation (SCA), WMMSE, and matching theory, have been employed for power allocation [122, 236, 237, 238, 239, 240]. However, a major drawback of these methods is their reliance on iterative algorithms, which are computationally intensive and result in long runtimes, ultimately degrading system energy efficiency. The deep unfolding approach mitigates this issue by reducing the number of iterations required to achieve a near-optimal solution, thereby enhancing computational efficiency and energy savings.
Wireless networks expand in terms of spatial coverage, but traditional DNN architectures, such as multi-layer perceptrons and CNNs, struggle to capture the inherent topology of these networks. To address this limitation, Chowdhury et al. [241] propose leveraging GNNs to parameterize the power allocation function. In their approach, the WMMSE algorithm is unfolded into layers, resulting in a distributed, efficient, and scalable solution. The unfolded WMMSE method integrates domain-specific elements with trainable components, which are parameterized by GNNs to exploit the natural graph structure of wireless communication networks. The objective is to map the channel matrix and approximate optimal power allocation by training the GNN using a block coordinate descent algorithm.
In [242], Li et al. unroll the SCA algorithm into NN layers for energy-efficient power allocation in wireless interference networks. The proposed unfolded algorithm is highly adaptable, as it generalizes across varying network sizes and channel distributions while also enabling transfer learning between different path loss models. Given that the system operates in a multi-user wireless interference network, it can be naturally modeled as a graph neural network GNN. The power allocation policy is formulated based on the CSI matrix ${\bf{H}}$ and the total power constraint $P_{\text{total}}$, within a layered architecture incorporating trainable weights.
@@
Hu et al. [243] propose an unsupervised power allocation optimization approach for ad-hoc networks by integrating attentive graph representation with deep unfolded WMMSE. The key advantage of attentive graph representation is its ability to allow network nodes to selectively focus on the most relevant connections during learning, rather than treating all links equally. The power allocation framework is built upon this attentive graph representation, as described in [244, 245]. Meanwhile, the deep unfolded WMMSE maps the iterative steps of the WMMSE algorithm into neural network layers within an unsupervised learning paradigm. This approach effectively combines deep unfolding with graph representation learning, significantly reducing the search space for the objective function. As a result, the model achieves improved interpretability and accelerated training while eliminating the need for supervised data.
X Physical Layer Security
Physical layer security is a critical research area aimed at ensuring secure and reliable transmission between the transmitter and receiver while preventing eavesdropping and impersonation attacks. Due to the broadcast nature of wireless communication, adversarial users pose a significant threat. These adversaries generally fall into two categories: eavesdroppers (or middleman attackers), who attempt to extract information from an ongoing transmission without detection, and jammers, who deliberately degrade the signal quality to prevent the receiver from successfully decoding the intended message [246]. Developing robust security mechanisms at the physical layer is essential to mitigate these threats and enhance communication reliability.
The fundamental network model for analyzing security and confidentiality is a three-terminal system, as illustrated in Fig. 8. In this setup, the legitimate transmitter (Alice), the intended receiver (Bob), and the unauthorized eavesdropper (Eve) are equipped with $N_{\text{t}}$, $N_{\text{r}}$ and $N_{\text{e}}$ antennas, respectively. As introduced by Wyner in his seminal work on wiretap channels [247], this model captures the imperfections introduced by the communication channel. Alice transmits an information signal, $\bf{x}$, a random vector intended for Bob, over the main channel, which is modeled as a discrete memoryless channel. However, due to the broadcast nature of wireless communication, the signal received by Bob, $\bf{y}$, is also partially intercepted by Eve through an additional wiretap channel, resulting in her received signal, $\bf{z}$, as shown in Fig.  8.
Figure 8: General wiretap channel [246].
Traditionally, secret key-based encryption has been the standard approach to ensuring message confidentiality, where the encryption key is shared exclusively between the transmitter and the intended receiver. As defined in [248], a secrecy system can be abstractly described as a set of transformations that map a set of possible messages to a set of possible cryptograms. Each transformation corresponds to an encryption process that utilizes a unique, secure key to encipher the message, ensuring that unauthorized entities cannot decipher the transmitted information.
The primary objective of a secure communication system is to maximize the secrecy rate, ensuring that the transmission rate to the intended user remains high while minimizing information leakage to the eavesdropper. The secrecy capacity $C_{s}$ is defined as the maximum rate $R_{s}$ at which secure communication can be achieved, guaranteeing that the eavesdropper receives negligible information about the transmitted message. This capacity represents the fundamental limit of secure data transmission, balancing efficiency and confidentiality in the presence of adversarial interception. The secrecy capacity can be expressed as:
$C_{s}=\max_{P_{X}}\left\{I\left({\bf{X};\bf{Y}}\right)-I\left({\bf{X};\bf{Z}}\right)\right\}$
where, $I(\bf{X};\bf{Y})$ is the mutual information between the transmitter and the intended user, and $I(\bf{X};\bf{Z})$ represents the information leakage to the eavesdropper. The transmitter can maximize the secrecy capacity by carefully choosing the probability distribution $P_{X}$ for the transmitted signal $\bf{x}$ to maximize the difference $I(\bf{X};\bf{Y})$-$I(\bf{X};\bf{Z})$ i.e, the rate $R_{s}$.

### v3

@@ VI-A MIMO ISAC
The large-scale antenna arrays in mmWave MIMO systems enable sub-meter user localization by providing high angular resolution. However, these methods often incur significant computational complexity, particularly in high dimensional mmWave MIMO settings. To improve this, Fan et al. proposed, a fast and accurate direct localization method in [193]. An ADMM solver is developed to solve the optimization problem with low complexity. The ADMM iterations are unfolded to learn the trainable parameters that include penalty and proximal parameters, and weight coefficient for channel gain between the BS and user. The unfolded network converges faster than classical ADMM and computational time is reduced.
VI-B Modern Multicarrier Modulation for ISAC
A joint passive sensing and communication demodulation for OFDM is addressed in [194]. The ISAC-Net is designed by unfolding two algorithms: a denoising-based AMP algorithm for the communication module and the 2D discrete Fourier transform (DFT) for the sensing module. The step-size to update the channel and DFT matrix form the trainable set to balance the communication demodulation and sensing accuracy. Furthermore, Jingcai et al. proposed an unfolded network (FISTA-Net) for joint target parameter estimation and communication in MIMO-OFDM systems [195]. The fast ISTA is unfolded, with learnable parameters such as step-size for weight matrix, enabling high-precision solutions with low computational complexity. The proposed network converges with $8$ layers as compared to $2000$ iteration in fast ISTA. Also, the network achieves an NMSE of $-20.12$ dB while the classical complex ISTA achieved $-16.8$ dB. The problem of estimating parameters of moving targets in high-speed scenarios using OTFS modulation for ISAC systems is explored in [196]. The ADMM iterations are unfolded (ADMM-Net) to accelerate convergence and enhance sparse recovery of target position. Regarding parameter estimation, ADMM-Net [196] and FISTA-Net [195] reduce NMSE significantly. Also, ISAC-Net [194] and FISTA-Net [195] show faster convergence than iterative baselines.
VI-C IRS-Assisted ISAC
Liu et al. addressed a robust beamforming design for an IRS-assisted ISAC system under imperfect CSI and target direction estimates in [197]. The objective is to minimize the sum of worst-case Cramér-Rao bounds (CRB) for all sensing targets. They formulated a bi-level optimization structure that consists of two levels: the lower-level problem finds the worst-case CSI and worst-case target sensing angles. The upper-level problem optimizes the active and passive beamforming matrices. A double-loop deep unfolding NN is designed to tackle the bi-level optimization problem. The CSI and angle uncertainties are updated via PGD in the lower-level loop. The beamforming variables are updated in the upper-level loop using PGD and Riemannian MO optimization. The proposed network effectively achieves near-optimal performance by handling statistical CSI errors and bounded target direction errors.
VI-D Discussions and Takeaways
Deep unfolding offers a structured and interpretable framework for tackling challenging optimization problems in ISAC systems. By unrolling classical algorithms like PGD and WMMSE into NN layers, unfolding enables the joint design of beamformers and waveforms that balance sensing accuracy and communication performance. These approaches reduce computational complexity and provide faster convergence compared to traditional iterative solutions. We summarize typical deep unfolding models for ISAC in Table VI. It is observed that deep unfolding exhibits advantages in constant modulus waveform design, optimization of dual-functional beamformers, and power allocation. Indeed, the deep unfolding technique can effectively handle non-convex constraints or objective functions like SINR, sum rate, and CRB, which are infeasible with black-box models.
In [198], the authors have provided techniques such as gradient calculation for estimating position that can be employed to facilitate practical deployment. However, the application of deep unfolding in ISAC is still underexplored. Indeed, most existing unfolding solutions are developed to optimize the transmit waveform in monostatic ISAC scenarios. In sensing functions, the processing of the received echo signal plays an important role in translating the observation into a decision on target detection and localization. Well-designed deep unfolding models can perform these tasks with low complexity and high accuracy, which is important for real-time sensing applications. There is limited literature addressing the practical deployment of unfolded networks in ISAC systems.
VII Decoding in Error Correcting Codes
Mobile communication has evolved across generations. 1G used analog transmission (without any coding), and 2G introduced digital voice coding with convolutional and block codes. Turbo codes, adopted in 3G and 4G, offer near-Shannon limit performance with moderate complexity [199]. Low-density parity check (LDPC) codes, rediscovered in the 1960s, are now widely used in 4G and 5G [200], [201], [202]. Polar codes, introduced in 2009, achieve the capacity of symmetric binary-input discrete memoryless channels [203], approaching the Shannon limit as the block length increases. They have been adopted in 5G new radio for both uplink and downlink control channels [204]. BP operates on a bipartite graph known as Tanner graph, introduced by Michael Tanner [205] The min-sum and max-sum algorithms are derived from the generic sum-product decoding [206]. Polar codes use successive cancellation decoding, where the received bits are decoded sequentially by leveraging frozen bits and previously decoded bits. More advanced techniques, such as successive list decoding assisted by cyclic redundancy check (CRC), further improve the performance of polar codes [207].
Wei et al. in [39] proposed an unfolded ADMM-based penalized algorithm for binary linear codes to enhance the decoding performance. The learnable ADMM decoding network (LADN) trains the ADMM penalty parameters. Additionally, decoding irregular binary LDPC codes via unfolded ADMM is explored in [40]. The LADN [39] outperforms the original ADMM-penalized decoder in block error rate for LDPC codes, while the network in [40] outperforms the original ADMM-penalized decoder in block error rate for irregular LDPC codes.
In [36], the authors proposed the TurboNet architecture, to improve turbo decoding by unfolding the max-log MAP algorithm. Trainable parameters were inserted into the max-log-MAP decoding flow like weights for prior LLRs. Numerical analysis demonstarted that TurboNet outperforms classical turbo decoding based on the max-log-MAP algorithm, particularly at moderate-to-high SNRs. Furthermore, TurboNet has been validated through over the air experiments confirming its robustness to SNR mismatches. Moreover, the authors have extended the work by including joint detection and decoding in [208]. A CGD-based OAMP detector (CG-OAMPNet) for the MIMO-OFDM system is proposed by unfolding EP. The robustness of CG-OAMPNet was further validated through over-the-air experiments, demonstrating practical resilience to environmental variations. In correlated channel scenario, CG-OAMPNet outperforms OAMP-Net2 [11] by performance gain of $\sim 2$ dB at BER$=10^{-3}$. The authors in [209] tackle the decoding challenges of marker codes in insertion and deletion channels without perfect CSI. They proposed a network called FBNet to improve detection performance; the forward-backward algorithm is unfolded for marker codes. The trainable weights are the CSI probabilities for insertion/deletion channels. FBNet performs quite close to the classical forward-backward algorithm but is robust to CSI mismatch. FBNet [209] is based on RNN structure and target synchronization errors (insertions/deletions) without accurate CSI, while TurboNet [36] is a deep feedforward DNN structure and targets substitution errors in turbo codes under AWGN channel.
Gao et al. [210] have unfolded the polar BP algorithm to improve polar code decoding performance under one-bit quantization. The network with trainable weights for the factor graph edges outperforms standard BP with two bit quantizer by $0.2$ dB. In [211], a double-loop iterative decoding algorithm for LDPC codes is unfolded within a penalty dual decomposition framework. The network alternately updates primal and dual variables in a double loop structure. The penalty coefficients, penalty parameters and over relaxation factor form the trainable parameter set. The network has lower computational complexity than ADMM based decoders. Additionally, by unfolding the iterative decoding process between check nodes and variable nodes for LDPC codes, a deep unfolded normalized min-sum LDPC decoder is implemented in [212]. The unfolded network outperforms the classical normalized min-sum LDPC decoder in BER and is robust across code lengths, rates and different channel conditions. The networks in [211, 212] focus on LDPC codes while the network in [210] consider polar codes. Moreover, the networks in [211, 212] use multi-SNR training for robustness.
@@ VII Decoding in Error Correcting Codes
In [213], theoretical results for the unfolded BP decoder for Tanner codes are presented. The authors define the generalization gap of a decoder as the difference between the empirical and expected BER. They provide a set of theoretical results that bound this gap and examine its dependence on decoder complexity, which is influenced by code parameters (such as block length, message length, and variable/check node degrees), the number of decoding iterations, and the size of the training dataset.
Traditionally, channel estimation, signal detection, and channel decoding in a MIMO receiver are performed sequentially and independently. This disjoint approach leads to suboptimal performance. To address this, Sun et al. [214] propose an unfolded ADMM-based receiver for joint detection and decoding. The detection and LDPC decoding components are jointly unfolded from ADMM iterations. Simulations show that the proposed receiver outperforms classical turbo receivers. In [215], Zhang et al. proposed a joint signal detection and channel decoding for MIMO systems. Signal detection is achieved by unfolding the EP algorithm. The channel decoding is based on an unfolded turbo code decoding network, TurboNet, as proposed in [36]. The damping factors in EP and decoder parameters in turbo decoding are learned in the training stage. The networks showed fast adaptation and strong BER gains in over-the-air experiments.
TABLE VII: Summary of typical deep unfolding models for decoding and their advantages and disadvantages.
Unfolding Model Main Novelties and Performance Gains Advantages and Disadvantages
LADN [39] Introduces learnable ADMM-based decoder with trainable penalties for binary LDPC decoding. Achieves a block error rate reduction of 10$\sim$ 10$\times$ 10 × at SNR $=2$ dB over the baseline ADMM decoder. Advantages: Has small number of learning parameters and fast convergence. Disadvantages: Limited by code length. Inherent codeword structure not exploited leading to high computational complexity.
TurboNet [36] Unfolds max-log MAP decoding steps for turbo codes with learnable weights. Outperforms the conventional max-log MAP scheme by $0.25-0.5$ dB BER. Advantages: Compatible with 3GPP turbo codes and validated through practical experiment. Disadvantages: Training requires fixed interleaver and careful SNR selection.
Quantized Polar BP [210] Learns BP decoding under 1-bit quantization for polar codes over an AWGN channel. Outperforms standard BP with one-bit quantization by $1.2-2$ dB (block error rate). Advantages: Hardware-friendly and applicable to IoT systems with coarse ADCs. Disadvantages: Less scalable. Require large number of trainable weights assigned to each edge of factor graph for decoding.
ADMM-Net [40] Extends LADN [39] to irregular LDPC codes. Unfolded ADMM, learn different penalty weights for each variable node in decoding. Achieve better BER performance than BP and ADMM penalized decoders for irregular LDPC codes. Advantages: Suitable for irregular code graphs. Disadvantages: The initialization strategy for the trainable parameters assumes prior knowledge of the code structure, which makes it less generalizable.
FBNet [209] Introduce a RNN-based unfolding model based on forward-backward decoding for Marker codes over insertion-deletion channels. Outperforms the classical forward-backward algorithm under imperfect CSI. Advantages: Robust to synchronization errors with flexible sequence handling. Disadvantages: Degraded performance under burst insertion-deletion channels.
VII-A Discussions and Takeaways
In Table VII, we summarize the most popular unfolded network for the decoding of channel codes. Deep unfolding spans their applications in modern error-correcting codes, such as LDPC, turbo, and polar codes. The soft probabilistic estimates are progressively refined through learnable decoding layers in deep unfolding while enforcing parity constraints, enabling near-optimal hard decisions without the need for manual hyperparameter tuning. Hence, it merges model based optimization, for structured soft-output refinement and codeword validity with data-driven learning, for adaptive penalty tuning, outperforming traditional decoders in accuracy and efficiency.
Despite these advantages, there are several limitations in existing deep unfolding architectures for decoding and channel coding. For example, the unfolded decoders in [39] is tied to specific code structures, requiring retraining when parameters such as code length or interleaving patterns change. Some methods demand significant training effort or hyperparameter tuning to maintain generalization.
There are several works [215, 36, 208] that developed a prototyping system to experimentally validate the effectiveness and practical feasibility of unfolded networks in real-world channel conditions. The experimental results indicated that the unfolded network demonstrated strong robustness to environmental variations and operate effectively without the need for frequent retraining.
Future research should aim to create more adaptable unfolding architectures that can handle a variety of code structures, channel models, and SNR conditions. Joint optimization with modulation and detection, as well as hardware-aware training, may further improve the viability of deep unfolding for end-to-end error correction in next-generation wireless communication systems.
VIII Power Allocation
@@
Power allocation in wireless communication refers to the process of distributing available transmit power across different communication channels or users to optimize system performance. It aims to maximize throughput, minimize interference, and ensure fairness, while considering factors like channel conditions, user requirements, and energy efficiency for improved overall network efficiency. Several classical approaches, such as the Dinkelbach algorithm, Lagrange dual decomposition, interference pricing, successive convex approximation (SCA), WMMSE, and matching theory, have been employed for power allocation [144, 216, 217, 218, 219, 220]. However, a major drawback of these methods is their reliance on iterative algorithms, which are computationally intensive and result in long runtimes, ultimately degrading system energy efficiency. The deep unfolding approach mitigates this issue by reducing the number of iterations required to achieve a near-optimal solution, thereby enhancing computational efficiency and energy savings.
@@
Chowdhury et al. in [221] unfolded WMMSE into GNN based architecture to learn a power allocation policy from the channel matrix. The key feature of their architecture is that graph topology is leveraged with GNN for efficient rate utility maximization. The unfolded WMMSE method integrates domain-specific elements with trainable components, which are parameterized by generated adaptive weigh of the GNN. Hu et al. [222] propose an unsupervised power allocation optimization approach for ad-hoc networks by integrating attentive graph representation with deep unfolded WMMSE. The key advantage of attentive graph representation is its ability to allow network nodes to selectively focus on the most relevant connections during learning, rather than treating all links equally. The power allocation framework is built upon this attentive graph representation, as described in [223, 224]. The model [222] achieves improved interpretability and accelerated training over [221] while eliminating the need for supervised data. In [225], Li et al. unroll the SCA algorithm into NN layers for energy-efficient power allocation in wireless interference networks. The proposed unfolded algorithm is highly adaptable, as it generalizes across varying network sizes and channel distributions. The power allocation policy is formulated based on the CSI matrix and the total power constraint, within a layered architecture incorporating trainable weights. Each of the networks discussed in [221, 222] and [225] leverages the GNN structure in the deep unfolding approach and achieve near optimal performance of their benchmark iterative algorithm i.e, WMMSE and SCA respectively.
IX Physical Layer Security
Physical layer security aims to ensure secure and reliable communication while preventing eavesdropping and impersonation attacks. Wireless communication’s broadcast nature exposes it to adversaries like eavesdroppers, who intercept transmissions, and jammers, who degrade signal quality. The primary objective is to maximize the secrecy rate, ensuring high transmission rates for the intended user while minimizing information leakage to the eavesdropper. The secrecy capacity represents the maximum rate at which secure communication can be achieved, balancing efficiency and confidentiality. It is defined as the difference between mutual information between transmit and receive signals and information leakage to the eavesdropper [226].